OpenStudy (anonymous):

1/2 + 4/2x = x+4/10 x=-4 x=5 x=-4 and x=5 no solution

OpenStudy (anonymous):

4/2 x = 2x , 1/2 = 5/10 1/2 + 4/2x = x + 4/10 5/10 + 2x = x + 4/10 subtract 5/10, subtract x x = -1/10 should be your answer

OpenStudy (anonymous):

so no solution then?

OpenStudy (anonymous):

well there IS a solution, it's x = -1/10 (try putting it into the equation, you will get 3/10 = 3/10 however none of the given answers are solutions, try them out yourself also.

OpenStudy (anonymous):

so i guess the answer would be "no solution" however it's incorrect lol in my opinion

OpenStudy (anonymous):

ok thanks. I had tried it out and didnt get a solution(for the given choices) but I just wanted to be sure!